OpenStudy (amina_soneviseth):

MEDALS ARE GIVEN how to find the inequity. 5 = - (-z+3) a.) -8 b.)8 c.)-2 d.)2 ------------------------------

OpenStudy (anonymous):

i think a or c

OpenStudy (amina_soneviseth):

okay thank you! :D

OpenStudy (anonymous):

sorry if that did not help .-.

OpenStudy (amina_soneviseth):

lol i know its eight but is it negative?

OpenStudy (anonymous):

Yes it's negative

OpenStudy (trojanpoem):

How can this be inequality ? Nevertheless, 1) Divide over -1 -5 = -z +3 2) Let 3 be in the other side -z = -8 3) remove the signs together z= 8

OpenStudy (jango_in_dtown):

5 = - (-z+3) gives 5=z-3 gives 5+3=z and hence z=8

OpenStudy (amina_soneviseth):

actually i did the problem alone and its positive lol

OpenStudy (anonymous):

good job!